Question 4 Solve the differential equation. 2xy + y = 2V*Question 5 Solve the initial value problem xy + y = xln x , y(1) = 0Question 7 Find the derivative. c = tet, g =t+ sin tQuestion 8 Find the equation of the tangent to the curve at the given point. x = ť – t, y=ť +t+1 ; (0,3)
Answer #1

14) solution Given differential equation is 2xy+y = 2 va Dividing on both side with ſal, we get yt my = on tehnyt -(0) Equy. Ta = fax tc y va = a + c y = a (a + c) y = a + c. all (5 solution Given dififerential equation is ay+y = almal Dividing oThe general solution of the given differential equation is Y (IF ) - Q.I.F 18+ Y.x = [X \na da +c 2y = [int]x dx. - Strony )put the value of c in equation@ 19 = 4 Ina - pop + 42 (7) solution - Given that a = tet and y=t+Sint di = (tet) = ct et 4 tiedt = (t2-1) = 27-1 (+2+7 + 1) = 2t + dy = 27+) da att c.. where t=1) day cut Co, 3) is m = 2015+) = 3 = 3 20171 The equation